# Function-Correcting Partition codes

### v1

@@ III-A FCPCs for protecting multiple functions
Solution 1: Use an error-correcting code for all $k=35$ symbols that can correct up to $t=2$ error for which the minimum required length is $46$ [17].
@@
Since $3\geq t+1$ for $t=2$, we have an optimal $(\Delta_{3},t)$-FCC with redundancy $2t=4$ which is an $(\mathcal{P}_{\Delta_{3}},t)$-encoding. Therefore, we can use this encoding by which the desired functions of both receivers will be protected from two errors, and the length will be $k+2t=35+4=39.$
The savings achieved by employing a single code to protect multiple partitions (or functions) can be quantified in two ways: (i) the saved redundancy per function, referred to as the partition redundancy gain; and (ii) the relative improvement in the code rate, referred to as the partition rate gain.
@@
Let $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{K}$ be partitions of $\mathbb{F}_{q}^{k}$, and let $\mathcal{P}=\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}$. The partition redundancy gain of an $(\mathcal{P},t)$-encoding with redundancy $r$ is defined as
@@
The partition rate gain, defined as the relative rate increment, is given by
@@
These partition gains can be seen in terms of functions also. Let $f^{(1)},f^{(2)},\ldots,f^{(K)}$ be functions on $\mathbb{F}_{q}^{k}$ with domain partitions $\mathcal{P}_{f^{(1)}},\mathcal{P}_{f^{(2)}},$ $\ldots,$ $\mathcal{P}_{f^{(K)}}$, respectively, and $\mathcal{P}=\mathcal{P}_{f^{(1)}}\vee\mathcal{P}_{f^{(2)}}\vee\cdots\vee\mathcal{P}_{f^{(K)}}$. Then the partition redundancy gain of an $(\mathcal{P},t)$-encoding with redundancy $r$ is
@@
In a similar manner, the partition rate gain is $\frac{\sum_{i=1}^{K}r_{f^{(i)}}(k,t)-r}{k+r}$.
For the code given in solution 3 of Example 3, the partition redundancy gain is $\frac{1}{2}((4+4)-4)=2,$ and the partition rate gain is $\frac{4}{39},$ which are the maximum possible gains as $(\mathcal{P}_{\Delta_{3}},2)$-encoding used here is optimal with $r_{\mathcal{P}_{\Delta_{3}}}(35,2)=4$.
@@
$\max_{i\in[K]}(r_{\mathcal{P}_{i}}(k,t))\leq r_{\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}}(k,t)\leq\min\left(N(q^{k},2t+1)-k,\sum_{i=1}^{K}r_{\mathcal{P}_{i}}(k,t)\right),$
where $N(M,d)$ denotes the minimum length of an error-correcting code with $M$ codewords and minimum distance $d$.
@@
The first part of the upper bound is quite straightforward, since any $(n,q^{k},2t+1)$ error-correcting code is an $(\mathcal{Q},t)$-encoding for the finest partition $\mathcal{Q}=\{\{x\}:x\in\mathbb{F}_{q}^{k}\}$, which is a refinement of any partition. Hence, by Lemma 2,
$r_{\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}}(k,t)\leq N(q^{k},2t+1)-k.$
@@
where $r_{i}(u)$ is redundancy vector of an $(\mathcal{P}_{i},t)$-encoding of length $r_{\mathcal{P}_{i}}(k,t)$.
@@
and from (1), we have $d(C(u),C(v))\geq 2t+1$. Therefore, $C$ is an $(\mathcal{P},t)$-encoding, and hence
@@ III-B FCPCs for linear functions
Let $U$ be a subspace of $\mathbb{F}_{q}^{k}$ and $\mathcal{U}$ denote the coset partition of $U$. Then for a positive integer $t$, $(\mathcal{U},t)$-encoding is an $(f,t)$-FCC for any linear function $f:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{\ell}$ with Ker(f)=U(f)=U.
@@
If $f^{(1)},f^{(2)},\ldots,f^{(K)}$ are linear functions from $\mathbb{F}_{q}^{k}$ to $\mathbb{F}_{q}^{\ell}$. Then a $(\mathcal{U},t)$-encoding is an $(f^{(i)},t)$-FCC for all $i\in[K]$, if $\mathcal{U}$ is the coset partition of subspace $U=\ker(f^{(1)})\cap\ker(f^{(2)})\cap\cdots\cap\ker(f^{(K)})$.
@@
for some $x_{i}\in\mathbb{F}_{q}^{k},i\in[K]$. Since $X\neq\emptyset$, we have $x\in X$. That means
@@
Consider two functions $f_{1},f_{2}:\mathbb{F}_{2}^{3}\to\{0,1\}$ defined as
@@
This systematic code works as both $(f_{1},1)$-FCC and $(f_{2},1)$-FCC. If we construct an FCC only for $f_{1}$, then we need only $2$ redundancy symbols: assign $00$ to all vectors in $\ker(f_{1})$ and $11$ to all vectors in $100+\ker(f_{1})$. This $(f_{1},1)$-FCC is optimal, i.e., $r_{f_{1}}(3,1)=2$. Similarly, we have $r_{f_{2}}(3,1)=2$. Therefore, the partition redundancy gain of $(\mathcal{U},t)$-encoding given above is $\frac{1}{2}(2+2-3)=\frac{1}{2}$, and the partition rate gain is $\frac{1}{6}$.
III-C Function-correcting partition code construction for Locally bounded partitions
@@
For $\lambda=2$, every Hamming ball of radius $\rho$ intersects at most two blocks of the partition $\mathcal{P}$.
We now specialize to $(2t,2)$-locally bounded partitions and present a $(\mathcal{P},t)$-encoding for it.

### v2

@@ III-A FCPCs for protecting multiple functions
Solution 1: Use an ECC for all $k=35$ symbols that can correct up to $t=2$ error for which the minimum required length is $46$ [18].
@@
Since $3\geq t+1$ for $t=2$, we have an optimal $(\Delta_{3},t)$-FCC with redundancy $2t=4$ which is a $(\mathcal{P}_{\Delta_{3}},t)$-encoding. Therefore, we can use this encoding by which the desired functions of both receivers will be protected from two errors, and the length will be $k+2t=35+4=39.$
The savings achieved by employing a single code to protect multiple partitions (or functions) can be quantified in two ways: (i) the saved redundancy per function, referred to as the partition redundancy gain; and (ii) the relative improvement in the code rate, referred to as the relative rate increment.
@@
Let $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{K}$ be partitions of $\mathbb{F}_{q}^{k}$, and let $\mathcal{P}=\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}$. The partition redundancy gain of a $(\mathcal{P},t)$-encoding with redundancy $r$ is defined as
@@
The relative rate increment is defined as
@@
These partition gains can be seen in terms of functions also. Let $f^{(1)},f^{(2)},\ldots,f^{(K)}$ be functions on $\mathbb{F}_{q}^{k}$ with domain partitions $\mathcal{P}_{f^{(1)}},\mathcal{P}_{f^{(2)}},$ $\ldots,$ $\mathcal{P}_{f^{(K)}}$, respectively, and $\mathcal{P}=\mathcal{P}_{f^{(1)}}\vee\mathcal{P}_{f^{(2)}}\vee\cdots\vee\mathcal{P}_{f^{(K)}}$. Then the partition redundancy gain of a $(\mathcal{P},t)$-encoding with redundancy $r$ is
@@
In a similar manner, the relative rate increment is $\frac{\sum_{i=1}^{K}r_{f^{(i)}}(k,t)-r}{k+r}$.
For the code given in solution 3 of Example 3, the partition redundancy gain is $\frac{1}{2}((4+4)-4)=2,$ and the relative rate increment is $\frac{4}{39},$ which are the maximum possible gains as $(\mathcal{P}_{\Delta_{3}},2)$-encoding used here is optimal with $r_{\mathcal{P}_{\Delta_{3}}}(35,2)=4$.
@@
$\max_{i\in[K]}(r_{\mathcal{P}_{i}}(k,t))\leq r_{\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}}(k,t)\leq\min\left(N_{q}(q^{k},2t+1)-k,\sum_{i=1}^{K}r_{\mathcal{P}_{i}}(k,t)\right),$
where $N_{q}(M,d)$ denotes the minimum length of an ECC with $M$ codewords and minimum distance $d$.
@@
The first part of the upper bound is quite straightforward, since any $(n,q^{k},2t+1)$ ECC is an $(\mathcal{Q},t)$-encoding for the finest partition $\mathcal{Q}=\{\{x\}:x\in\mathbb{F}_{q}^{k}\}$, which is a refinement of any partition. Hence, by Lemma 2,
$r_{\mathcal{P}_{1}\vee\mathcal{P}_{2}\vee\cdots\vee\mathcal{P}_{K}}(k,t)\leq N_{q}(q^{k},2t+1)-k.$
@@
where $r_{i}(u)$ is the redundancy vector of a $(\mathcal{P}_{i},t)$-encoding of length $r_{\mathcal{P}_{i}}(k,t)$.
@@
and from (1), we have $d(C(u),C(v))\geq 2t+1$. Therefore, $C$ is a $(\mathcal{P},t)$-encoding, and hence
@@ III-B FCPCs for linear functions
Let $U$ be a subspace of $\mathbb{F}_{q}^{k}$ and $\mathcal{U}$ denote the coset partition of $U$. Then for a positive integer $t$, an $(\mathcal{U},t)$-encoding is an $(f,t)$-FCC for any linear function $f:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{\ell}$ with Ker(f)=U(f)=U.
@@
If $f^{(1)},f^{(2)},\ldots,f^{(K)}$ are linear functions from $\mathbb{F}_{q}^{k}$ to $\mathbb{F}_{q}^{\ell}$. Then a $(\mathcal{U},t)$-encoding is an $(f^{(i)},t)$-FCC for all $i\in[K]$, if $\mathcal{U}$ is the coset partition of the subspace $U=\ker(f^{(1)})\cap\ker(f^{(2)})\cap\cdots\cap\ker(f^{(K)})$.
@@
for some $x_{i}\in\mathbb{F}_{q}^{k},i\in[K]$. Since $X\neq\emptyset$, we have $x\in X$. This means
@@
Consider the two functions $f_{1},f_{2}:\mathbb{F}_{2}^{3}\to\{0,1\}$ defined as
@@
This systematic code works as both $(f_{1},1)$-FCC and $(f_{2},1)$-FCC. If we construct an FCC only for $f_{1}$, then we need only $2$ redundancy symbols: assign $00$ to all vectors in $\ker(f_{1})$ and $11$ to all vectors in $100+\ker(f_{1})$. This $(f_{1},1)$-FCC is optimal, i.e., $r_{f_{1}}(3,1)=2$. Similarly, we have $r_{f_{2}}(3,1)=2$. Therefore, the partition redundancy gain of $(\mathcal{U},t)$-encoding given above is $\frac{1}{2}(2+2-3)=\frac{1}{2}$, and the relative rate increment is $\frac{1}{6}$.
III-C FCPC construction for locally bounded partitions
@@
For $\lambda=2$, every Hamming ball of radius $\rho$ intersects at most two blocks of the partition $\mathcal{P}$. We now specialize to $(2t,2)$-locally bounded partitions and present a $(\mathcal{P},t)$-encoding for it.